OpenStudy (anonymous):

A. write four equivalent equations to x=8 using one of the four operations of addition, subtractions, division, multiplcation for each equivalent equation. use each operation only once. B. write an equivalent equation to x=8 that has the variable x on both sides. C.determine if n/6=15/18 and 2(n+1)=(n-1) are equivalent equations. determine if either equations is equivalent to any of the equations created for part A and B

OpenStudy (ash2326):

@levibutterfly25 . I'll do one part of A, say addition. You'll have to try subtraction, division and multiplication. Will you do that?

OpenStudy (anonymous):

sure

OpenStudy (ash2326):

We want to write equivalent expression to \[\large x=8\] So whatever operation we do, it should be done to both the sides. I'll do addition, I'll add my favorite digit 8 to both the sides \[\large x+8=8+8\] now we get \[\large x+8=16\] Now you try for subtraction, multiplication and division. If you get stuck, do let me know. I'll guide you:)